I am trying to solve and write an arithmetic equation for;
0,7,26,63,124,215. I have worked down to a common number (6). I don't know where to go from there.
Thank you for your help.
Emma

Answer by stanbon(75887)

0,7,26,63,124,215
------------
0
7 = 2^3-1
26 = 3^3-1
64 = 4^3-1
124 = 5^3-1
215 = 6^3-1
---
etc
Note: I have my doubts that the sequence is arithmetic.
Cheers,
Stan H.
==========